# Dents in the Mirror: A Novel Probe of Dark Matter Substructure in Galaxy Clusters from the Astrometric Asymmetry of Lensed Arcs

### v1

@@ 5.2 Measuring Asymmetry in Observed Arcs
First, we assume that in both arcs, the brightest pixel of each knot corresponds to its image position. Thus, we implicitly are ignoring transient and microlensing effects, and are assuming a very simple surface brightness for the source. This is likely a fair assumption for AS1063 System 1; however, for the Warhol arc this is a simplification. The Warhol arc has known lensed transients (yan23), some of which are likely to be contributing to the flux of the identified knots. Thus, the treatment of the transients in Warhol is ignored for now, and left for a future analysis of this arc.
@@ 5.3 Results
Figure 11: Posterior constraints on $f_{\rm sub}$ derived with our statistical method from AS1063 System 1 (solid black line) and the Warhol arc (dashed black line). The constraints from the Warhol arc should be treated as an upper limit. The solid red line indicates the combined posterior from the two arcs assuming the asymmetry detection in Warhol is genuine. The vertical dashed blue line indicates the combined posterior median, while the shaded blue region indicates the combined posterior’s 68% confidence interval. If we combine the two arcs, we constrain $\log f_{\rm sub}=-2.68^{+0.58}_{-0.74}$. We emphasize that these results are tentative, especially for AS1063 System 1, and elaborate on the limitations further in Section 5.3.
@@
Beginning with AS1063 System 1, we constrain at 68% CI $\log f_{\rm sub}=-2.36^{+0.56}_{-0.89}$. This is the first constraint on $f_{\rm sub}$ using our presented methodology. That being said, it is important to scrutinize the limitations of this constraint. The primary limitation concerns the underlying assumption that asymmetry is solely caused by astrometric perturbations from dark matter subhalos. It is important to realize that AS1063 System 1 is a very large perpendicular arc extending $\sim$7" perpendicularly from the critical curve, which is roughly twice the length of the fiducial perpendicular arc that we consider earlier in the paper. Therefore, it is less clear whether subhalos will be the dominant perturbing effect for the arc. Due to its size, it is possible that macrolens scale perturbations may contribute more to the arc’s asymmetry. To test this further, future studies should make use of multiple macrolens density models, rather than the single model (B19) that we use here. Nonetheless, our constraint is consistent with that found in previous studies (vegetti14; despali17; hsueh20).
For the Warhol arc, we infer at 68% CI the upper limit of $\log f_{\rm sub}$ to be $3.40^{+1.06}_{-0.97}$. As we saw with mock arcs, the relatively low asymmetry of this arc cause $f_{\rm sub}$ to be significantly less constrained.
Lastly, we implicitly assume that $f_{\rm sub}$ in both clusters results from the same underlying physics. As discussed in Section 3.5, we assume that both clusters have the same radial distribution of subhalos, have approximately the same mass, and that we are probing the cluster profile the same value of $R/R_{s}$. These assumptions are sufficient to first order based on CLASH measurements (umetsu14). This allows us to combine the two posteriors to achieve a joint constraint on $f_{\rm sub}$, much like in Section 4.4. We show this combined posterior in Figure 11. We find that $\log f_{\rm sub}=-2.68^{+0.58}_{-0.74}$. Future studies with our method will deliver more precise inferences from a greater number of arcs.
@@ 6 Conclusions
  • We apply our method to two real arcs. We make use of recent parametric lens models from LensTool for this exercise, and remeasure the arc image positions. We note that for this exercise, we make several simplifications and thus view our results as demonstrations of our method as applied to real data, rather than strict constraints on $f_{\rm sub}$. It is necessary to conduct more careful follow-up analyses for these two systems in the future. Even though our results are preliminary, it is still interesting to compare them to other results in the literature. For AS1063 System 1, we constrain $\log f_{\rm sub}=-2.36^{+0.56}_{-0.89}$ at 68% CI, which is consistent with previous constraints (vegetti14; despali17; hsueh20). For the Warhol arc, we constrain the upper limit of $\log f_{\rm sub}$ to be $3.40^{+1.06}_{-0.97}$ at 68% CI. We find the combined constraint to be $\log f_{\rm sub}=-2.68^{+0.58}_{-0.74}$ at 68%. We expect that future studies from a larger sample of lensed arcs will yield reliable, tight constraints on $f_{\rm sub}$.
There are many different directions of future work that we plan to pursue using this method. The easiest is to apply our method to a larger sample of lensed arcs, with the goal of obtaining a tight constraint on $f_{\rm sub}$. This work outlines this process using two well observed and modeled cluster lenses that allow for a quick application of our method. With recent high quality observations of cluster lenses from JWST, some cluster lenses have only very recently been updated in their lens models (e.g. the lenses observed in the Strong LensIng and Cluster Evolution (SLICE) program (cerny25)). Some candidate lensed arcs that appear to be usable with our method are the Quyllur arc in El Gordo (diego23c), Abell 68 System 1 (cerny25), Abell 2744 Systems 65 and 77 (furtak23), AS1063 System 2 (diego16b), SMACS J0723.3–7327 Systems 5 and 7 (mahler23b), and the $z\sim 10$ Cosmic Gems arc (bradley25). This is not an exhaustive list, and many others will be discovered with future observations allowing for higher precision identifications of lensed knots in arcs.
@@
Furthermore, we only consider a single lens macromodel for our analysis of AS1063 System 1 and the Warhol arc. A necessary improvement in future works will be to repeat our method for an ensemble of lens models. Combining the posteriors from the ensemble will also yield tighter constraints on $f_{\rm sub}$ for each arc. This is especially suitable for the Warhol arc, which has been the subject of recent interest in lens modelling with a variety of methods, including parametric (rihtarsic25; limousin25), free-form (cha23; perera24b), and hybrid (diego23a; cha25). Additionally, it was recently shown that lens models of MACSJ0416 continue to exhibit wide variability in mass reconstruction, despite the large increase in image constraints (perera25). To address this, it is necessary to marginalize our method over the variety of lens models available for the Warhol arc in order to eliminate bias from individual macrolens model assumptions. As we showed in this work, our method is resistant to macrolens profile, although somewhat different inferences will be made depending on the lens model’s reconstruction of the arc. For all these reasons, we recommend future studies to reexamine the Warhol arc and apply its myriad lens models with our method.
@@
An immediate extension of our analysis in the context of CDM is the inclusion of line-of-sight halos. Another direction of future work is to apply our method to different alternative models of dark matter. In this work, we only consider standard cold dark matter subhalos. The only changes in our method when using a different dark matter model would be the subhalo density profile (which would need to be re-calibrated to the respective dark matter model properties), and the subhalo mass function. In the warm dark matter paradigm, the primary changes would be the inclusion of free-streaming effects in the calculation of the tidal evolution, and the addition of a power-law term in the SHMF that includes the half-mode mass (lovell20). Applying our method for cluster lenses to constrain the half-mode mass of warm dark matter would compliment previous galaxy-scale lens constraints (gilman20). In the self-interacting dark matter (SIDM) framework, stringent upper limits on self-interaction cross section from cluster-scale lenses have been placed (jauzac16; andrade22). SIDM halos are more cored than standard CDM halos (nadler23) and have been modeled analytically (hou25) or as cored NFWs (gilman21; gilman23). The SHMF can be re-parameterized based on the fraction of subhalos undergoing core-collapse as a function of mass (gilman23). Like with warm dark matter, using our method to constrain the SIDM interaction cross section would compliment recent galaxy-scale constraints (e.g. kong24; tajalli25). For wave dark matter, constraints on the axion mass from cluster lenses are relatively sparse, although there have been some recent proposed methods (broadhurst25). Wave dark matter is distinct from warm dark matter and SIDM in that it can be modeled a Gaussian random field whose fluctuations are related to the axion mass (schive14). Reworking our method to constrain the wave dark matter axion mass would be one of the first constraints at cluster-scales.
@@ Appendix A Testing the Assumption of a Linear Critical Curve
Throughout this paper, we assume that on small scales for a smooth cluster-scale lens model, the critical curve is approximately linear. This assumption is motivated by previous analyses (venumadhav17; dai18) and is crucial to our analysis since the image midpoints will form along a straight line. It is important to determine the limits of this assumption, as this will determine whether it is appropriate to apply this assumption to observed arcs. This Appendix conducts two main tests: (1) A test of the limits of $\xi$ for deviations from a straight critical curve, and (2) A test of how much a deviation from linearity causes a bias in the inference of $f_{\rm sub}$. In the second test, we intend to show that any potential bias in the inference as a result of a nonlinear critical curve occur for arcs far more nonlinear than any considered in this work.
@@
Lastly, we test the effect of cluster member galaxies on the linear assumption. We note that cluster members that form very near to the arc will obviously distort the critical curve and induce asymmetries (e.g. diego23c). Instead of these obvious cases, we instead want to test the subtle presence of cluster members and how much they contribute to deviations from linearity. To test this, we add a small cluster member galaxy $\sim 3$" from the fiducial arc, as shown in Figure 13. The galaxy is located just outside of the subhalo window, such that it its presence is still noticeable in the background density profile, but not dominating the local mass distribution. As can be seen, the galaxy produces its own microcaustic near to the source position. The cluster member is modeled with a Singular Isothermal Sphere (SIS):

### v2

@@ 5.2 Measuring Asymmetry in Observed Arcs
First, we assume that in both arcs, the brightest pixel of each knot corresponds to its image position. Thus, we implicitly are ignoring transient and microlensing effects, and are assuming a very simple surface brightness for the source. This is likely a fair assumption for AS1063 System 1; however, for the Warhol arc this is a simplification. The Warhol arc has known lensed transients (Yan et al., 2023), some of which are likely to be contributing to the flux of the identified knots. Thus, the treatment of the transients in Warhol is ignored for now, and left for a future analysis of this arc.
@@ 5.3 Results
Figure 11: Posterior constraints on $f_{\rm sub}$ derived with our statistical method from AS1063 System 1 (solid black line) and the Warhol arc (dashed black line). The constraints from the Warhol arc should be treated as an upper limit. The solid red line indicates the combined posterior from the two arcs assuming the asymmetry detection in Warhol is genuine. The vertical dashed blue line indicates the combined posterior median, while the shaded blue region indicates the combined posterior’s 68% confidence interval. If we combine the two arcs, we constrain $\log f_{\rm sub}=-2.76^{+0.55}_{-0.72}$. We emphasize that these results are tentative, especially for AS1063 System 1, and elaborate on the limitations further in Section 5.3.
@@
Beginning with AS1063 System 1, we constrain at 68% CI $\log f_{\rm sub}=-2.44^{+0.61}_{-0.86}$. This is the first constraint on $f_{\rm sub}$ using our presented methodology. That being said, it is important to scrutinize the limitations of this constraint. The primary limitation concerns the underlying assumption that asymmetry is solely caused by astrometric perturbations from dark matter subhalos. It is important to realize that AS1063 System 1 is a very large perpendicular arc extending $\sim$7" perpendicularly from the critical curve, which is roughly twice the length of the fiducial perpendicular arc that we consider earlier in the paper. Therefore, it is less clear whether subhalos will be the dominant perturbing effect for the arc. Due to its size, it is possible that macrolens scale perturbations may contribute more to the arc’s asymmetry. To test this further, future studies should make use of multiple macrolens density models, rather than the single model (B19) that we use here. Furthermore, the large extent of this arc may no longer be suitable for the approximations of the image midpoints lying exactly on the critical curve (equation 17). We examine this possibility in greater detail in Appendix B. Nonetheless, our constraint is consistent with that found in previous studies (Vegetti et al., 2014; Despali and Vegetti, 2017; Hsueh et al., 2020).
For the Warhol arc, we infer at 68% CI the upper limit of $\log f_{\rm sub}$ to be $3.48^{+1.00}_{-0.91}$. As we saw with mock arcs, the relatively low asymmetry of this arc cause $f_{\rm sub}$ to be significantly less constrained.
Lastly, we implicitly assume that $f_{\rm sub}$ in both clusters results from the same underlying physics. As discussed in Section 3.5, we assume that both clusters have the same radial distribution of subhalos, have approximately the same mass, and that we are probing the cluster profile the same value of $R/R_{s}$. These assumptions are sufficient to first order based on CLASH measurements (Umetsu et al., 2014). This allows us to combine the two posteriors to achieve a joint constraint on $f_{\rm sub}$, much like in Section 4.4. We show this combined posterior in Figure 11. We find that $\log f_{\rm sub}=-2.76^{+0.55}_{-0.72}$. Future studies with our method will deliver more precise inferences from a greater number of arcs.
@@ 6 Conclusions
  • We apply our method to two real arcs. We make use of recent parametric lens models from LensTool for this exercise, and remeasure the arc image positions. We note that for this exercise, we make several simplifications and thus view our results as demonstrations of our method as applied to real data, rather than strict constraints on $f_{\rm sub}$. It is necessary to conduct more careful follow-up analyses for these two systems in the future. Even though our results are preliminary, it is still interesting to compare them to other results in the literature. For AS1063 System 1, we constrain $\log f_{\rm sub}=-2.44^{+0.61}_{-0.86}$ at 68% CI, which is consistent with previous constraints (Vegetti et al., 2014; Despali and Vegetti, 2017; Hsueh et al., 2020). For the Warhol arc, we constrain the upper limit of $\log f_{\rm sub}$ to be $3.48^{+1.00}_{-0.91}$ at 68% CI. We find the combined constraint to be $\log f_{\rm sub}=-2.76^{+0.55}_{-0.72}$ at 68%. We expect that future studies from a larger sample of lensed arcs will yield reliable, tight constraints on $f_{\rm sub}$.
There are many different directions of future work that we plan to pursue using this method. The easiest is to apply our method to a larger sample of lensed arcs, with the goal of obtaining a tight constraint on $f_{\rm sub}$. This work outlines this process using two well observed and modeled cluster lenses that allow for a quick application of our method. With recent high quality observations of cluster lenses from JWST, some cluster lenses have only very recently been updated in their lens models (e.g. the lenses observed in the Strong LensIng and Cluster Evolution (SLICE) program (Cerny et al., 2025)). Some candidate lensed arcs that appear to be usable with our method are the Quyllur arc in El Gordo (Diego et al., 2023), Abell 68 System 1 (Cerny et al., 2025), Abell 2744 Systems 65 and 77 (Furtak et al., 2023), AS1063 System 2 (Diego et al., 2016), SMACS J0723.3–7327 Systems 5 and 7 (Mahler et al., 2023), and the $z\sim 10$ Cosmic Gems arc (Bradley et al., 2025). This is not an exhaustive list, and many others will be discovered with future observations allowing for higher precision identifications of lensed knots in arcs.
@@
Furthermore, we only consider a single lens macromodel for our analysis of AS1063 System 1 and the Warhol arc. A necessary improvement in future works will be to repeat our method for an ensemble of lens models. Combining the posteriors from the ensemble will also yield tighter constraints on $f_{\rm sub}$ for each arc. This is especially suitable for the Warhol arc, which has been the subject of recent interest in lens modelling with a variety of methods, including parametric (Rihtaršič et al., 2025; Limousin et al., 2025), free-form (Cha and Jee, 2023; Perera et al., 2025b), and hybrid (Diego et al., 2024; Cha and Jee, 2026). Additionally, it was recently shown that lens models of MACSJ0416 continue to exhibit wide variability in mass reconstruction, despite the large increase in image constraints (Perera et al., 2025a). To address this, it is necessary to marginalize our method over the variety of lens models available for the Warhol arc in order to eliminate bias from individual macrolens model assumptions. As we showed in this work, our method is resistant to macrolens profile, although somewhat different inferences will be made depending on the lens model’s reconstruction of the arc. For all these reasons, we recommend future studies to reexamine the Warhol arc and apply its myriad lens models with our method.
@@
An immediate extension of our analysis in the context of CDM is the inclusion of line-of-sight halos. Another direction of future work is to apply our method to different alternative models of dark matter. In this work, we only consider standard cold dark matter subhalos. The only changes in our method when using a different dark matter model would be the subhalo density profile (which would need to be re-calibrated to the respective dark matter model properties), and the subhalo mass function. In the warm dark matter paradigm, the primary changes would be the inclusion of free-streaming effects in the calculation of the tidal evolution, and the addition of a power-law term in the SHMF that includes the half-mode mass (Lovell, 2020). Applying our method for cluster lenses to constrain the half-mode mass of warm dark matter would compliment previous galaxy-scale lens constraints (Gilman et al., 2020). In the self-interacting dark matter (SIDM) framework, stringent upper limits on self-interaction cross section from cluster-scale lenses have been placed (Jauzac et al., 2016; Andrade et al., 2022). SIDM halos are more cored than standard CDM halos (Nadler et al., 2023) and have been modeled analytically (Hou et al., 2025) or as cored NFWs (Gilman et al., 2021, 2023). The SHMF can be re-parameterized based on the fraction of subhalos undergoing core-collapse as a function of mass (Gilman et al., 2023). Like with warm dark matter, using our method to constrain the SIDM interaction cross section would compliment recent galaxy-scale constraints (e.g. Kong et al., 2024; Tajalli et al., 2025). For wave dark matter, constraints on the axion mass from cluster lenses are relatively sparse, although there have been some recent proposed methods (Broadhurst et al., 2025). Wave dark matter is distinct from warm dark matter and SIDM in that it can be modeled a Gaussian random field whose fluctuations are related to the axion mass (Schive et al., 2014). Reworking our method to constrain the wave dark matter axion mass would be one of the first constraints at cluster-scales.
@@ Appendix A Testing the Assumption of a Linear Critical Curve
Throughout this paper, we assume that on small scales for a smooth cluster-scale lens model, the critical curve is approximately linear. This assumption is motivated by previous analyses (Venumadhav et al., 2017; Dai et al., 2018) and is crucial to our analysis since the image midpoints will form along a straight line. It is important to determine the limits of this assumption, as this will determine whether it is appropriate to apply this assumption to observed arcs. This Appendix conducts two main tests: (1) A test of the limits of $\xi$ for deviations from a straight critical curve, and (2) A test of how much a deviation from linearity causes a bias in the inference of $f_{\rm sub}$. In the second test, we intend to show that any potential bias in the inference as a result of a nonlinear critical curve occur for arcs far more nonlinear than any considered in this work.
@@
Lastly, we test the effect of cluster member galaxies on the linear assumption. We note that cluster members that form very near to the arc will obviously distort the critical curve and induce asymmetries (e.g. Diego et al., 2023). Instead of these obvious cases, we instead want to test the subtle presence of cluster members and how much they contribute to deviations from linearity. To test this, we add a small cluster member galaxy $\sim 3$" from the fiducial arc, as shown in Figure 13. The galaxy is located just outside of the subhalo window, such that it its presence is still noticeable in the background density profile, but not dominating the local mass distribution. As can be seen, the galaxy produces its own microcaustic near to the source position. The cluster member is modeled with a Singular Isothermal Sphere (SIS):
@@
Appendix B Higher Order Lensing Effects Near the Critical Curve
In Section 2.2 we demonstrated that sources near the caustic fold will produce images that are symmetric across the critical curve (equation 17), and thus contain image midpoints that trace the critical curve. This result is valid in the limit that images are very close to the critical curve. It is therefore necessary to examine at what point this approximation breaks down, as larger arcs that contain knots further away from the critical curve may no longer contain midpoints along the critical curve. We examine this limit analytically, beginning first by expanding equation 13 to the third order:
$\boldsymbol{\beta_{i}}\left(\boldsymbol{\theta_{c}}+\delta\boldsymbol{\theta}\right)=\boldsymbol{\beta_{i}}(\boldsymbol{\theta_{c}})+\sum_{j}\frac{\partial\boldsymbol{\beta_{i}}}{\partial\boldsymbol{\theta_{j}}}\bigg\rvert_{\boldsymbol{\theta_{c}}}\delta\boldsymbol{\theta_{j}}+\frac{1}{2}\sum_{j,k}\frac{\partial^{2}\boldsymbol{\beta_{i}}}{\partial\boldsymbol{\theta}_{j}\partial\boldsymbol{\theta}_{k}}\bigg\rvert_{\boldsymbol{\theta_{c}}}\delta\boldsymbol{\theta_{j}}\delta\boldsymbol{\theta_{k}}\\ +\frac{1}{6}\sum_{j,k,l}\frac{\partial^{3}\boldsymbol{\beta_{i}}}{\partial\boldsymbol{\theta}_{j}\partial\boldsymbol{\theta}_{k}\partial\boldsymbol{\theta}_{l}}\bigg\rvert_{\boldsymbol{\theta_{c}}}\delta\boldsymbol{\theta_{j}}\delta\boldsymbol{\theta_{k}}\delta\boldsymbol{\theta_{l}}$ (31)
To solve, we proceed the same as we do in the Section 2.2, writing to leading order:
$\beta_{2}=\frac{1}{2}\frac{\partial^{2}\beta_{2}}{\partial\theta_{2}^{2}}\bigg\rvert_{0}\delta\theta_{2}^{2}+\frac{1}{6}\frac{\partial^{3}\beta_{2}}{\partial\theta_{2}^{3}}\bigg\rvert_{0}\delta\theta_{2}^{3}$ (32)
where this equation is the higher order equivalent of equation 15.
The goal is to solve equation 32 to determine the new form of equation 17, which we call $\delta\theta_{2}^{\rm tot}$. Since we are interested in how the symmetry is broken, we define: $\delta\theta_{2}^{\rm tot}=\delta\theta^{\rm init}_{2}+\delta\theta^{{}^{\prime}}_{2}$ where $\delta\theta^{\rm init}_{2}$ is the initial result from equation 17. This reduces the problem to solving for $\delta\theta^{{}^{\prime}}_{2}$, which is the correction term from including the third order expansion. To simplify notation, let $b=\frac{\partial^{2}\beta_{2}}{\partial\theta_{2}^{2}}\bigg\rvert_{0}$ and $c=\frac{\partial^{3}\beta_{2}}{\partial\theta_{2}^{3}}\bigg\rvert_{0}$, where it is implicitly assumed that $b>>c$. This is likely a good assumption for smooth potential variations, like those considered here, but would break down for small compact variations, like those from subhalos.
Plugging $\delta\theta_{2}^{\rm tot}$ into the third order expansion gives:
$\beta_{2}=\frac{1}{6}c\left(\delta\theta^{\rm init}_{2}+\delta\theta^{{}^{\prime}}_{2}\right)^{3}+\frac{1}{2}b\left(\delta\theta^{\rm init}_{2}+\delta\theta^{{}^{\prime}}_{2}\right)^{2}$ (33)
Since $\delta\theta^{{}^{\prime}}_{2}$ is very small, we continue by only keeping leading order terms of $\delta\theta^{\rm init}_{2}$ and $\delta\theta^{{}^{\prime}}_{2}$. We are assuming $\delta\theta^{\rm init}_{2}>>\delta\theta^{{}^{\prime}}_{2}$, such that $\beta_{2}$ remains approximately equal to the second order result (equation 15 in Section 2.2), allowing $\beta_{2}$ to cancel with the $\frac{1}{2}b\left(\delta\theta^{\rm init}_{2}\right)^{2}$ term:
$\frac{1}{6}c\left(\delta\theta^{\rm init}_{2}\right)^{3}+\frac{1}{2}c\left(\delta\theta^{\rm init}_{2}\right)^{2}\delta\theta^{{}^{\prime}}_{2}+b\delta\theta^{\rm init}_{2}\delta\theta^{{}^{\prime}}_{2}\approx 0$ (34)
Solving this with our approximations gives (with the denominator dominated by the term including $b$):
$\delta\theta^{{}^{\prime}}_{2}=-\frac{\frac{1}{6}c\left(\delta\theta^{\rm init}_{2}\right)^{2}}{\frac{1}{2}c\delta\theta^{\rm init}_{2}+b}\approx-\frac{c}{6b}\left(\delta\theta^{\rm init}_{2}\right)^{2}$ (35)
This result for $\delta\theta^{{}^{\prime}}_{2}$ is interpreted as the midpoint shift off the critical curve position. Importantly, the shift scales quadratically with the image distance from the critical curve $\delta\theta^{\rm init}_{2}$, meaning that this midpoint shift becomes important for increasingly larger distances from the critical curve.
In order to determine at what distance this midpoint shift becomes important, it is necessary to calculate the magnitude of $\delta\theta^{{}^{\prime}}_{2}$ for a cluster scale lens. This is analytically difficult as the derivative terms $b$ and $c$ cannot be solved generally. To simplify this problem analytically such that we can get a rough scale of the distances at which the midpoint shift becomes important, we consider the following setup.
We assume that the cluster lens can be approximated as an SIS with a power-law lens potential: $\psi(r)\propto r^{\alpha}$. This allows the derivative terms to be solved analytically because the critical curve is a circle located at the Einstein radius $\theta_{E}$. This is a subtly important simplification because it allows the critical direction to be radial. If circular symmetry is not present, a change of variable is required, complicating the problem. We note that the usage of an SIS is an imperfect approximation because an SIS will not form an image pair that straddles the critical curve. However, since the SIS is only used in this calculation to calculate the derivative terms analytically (i.e. the SIS is not used to calculate the image positions $\delta\theta^{\rm init}_{2}$) and scale $\delta\theta^{{}^{\prime}}_{2}$ with some property of the lens (in this case $\theta_{E}$), we justify this choice as sufficient.
With our simplifications and justifications stated, an SIS power-law lens potential gives:
$|\delta\theta^{{}^{\prime}}_{2}|\approx\frac{\alpha-3}{6\theta_{E}}\left(\delta\theta^{\rm init}_{2}\right)^{2}\propto\frac{\left(\delta\theta^{\rm init}_{2}\right)^{2}}{\theta_{E}}$ (36)
For an exact SIS, $\alpha=1$, so $|\delta\theta^{{}^{\prime}}_{2}|\approx(\theta_{E})^{-1}{\left(\delta\theta^{\rm init}_{2}\right)^{2}}/3$. To apply this as an approximation to the observed arcs all we need to do is calculate $\theta_{E}$. This result is intuitive, as the mass of the lens increases the shift in the midpoint position off the critical curve is reduced. For the fiducial perpendicular arc we examine in Section 3.1.2, this calculation predicts $\delta\theta^{{}^{\prime}}_{2}=0.02$". We measure the exact shift to be 0.04", which is reasonably close in spite of the simplifications. The disagreement is due to combination of simplifying assumptions, source displacement from the exact fold position, and complexity in the lens potential. Nonetheless, this gives a sense of the scale of the midpoint shift, where it is on the same order as the astrometric uncertainty. Importantly, relative to $\delta\theta^{\rm init}_{2}$ for this arc, which is observed to be 1.33”, $\delta\theta^{{}^{\prime}}_{2}/\delta\theta^{\rm init}_{2}=0.03$, making up a very small shift relative to the size of the arc.
For the real arcs we study in Section 5, we can repeat the calculation with greater interest since these arcs are much larger than the fiducial ones. For Warhol and AS1063 Sys 1, we calculate $\delta\theta^{{}^{\prime}}_{2}$ to be 0.03" and 0.1", respectively. Likewise, the relative shift to the arc size $\delta\theta^{{}^{\prime}}_{2}/\delta\theta^{\rm init}_{2}$ is 0.02 and 0.03 for Warhol and AS1063 Sys 1, respectively, once again indicating a small shift relative to the arc size. AS1063 Sys 1 is of greater interest since the shift is expected to be larger than the astrometric uncertainty, corroborating the complications previously discussed in Section 5.3.
We conclude that the midpoint shift off the critical curve is primarily due to a third order effect. The magnitude of the shift scales with the image distance from the critical curve and inversely scales with the mass of the lens. This effect can be captured in the astrometric uncertainty for our fiducial arcs, allowing us to conclude that it is a negligible effect to the efficacy of our method. In practice, however, greater care must be taken to acknowledge this midpoint shift, as evidenced by AS1063 Sys 1. We also emphasize that the shape of the lens potential can contribute to a highly nontrivial expected shift. Likewise, a constant offset could be a result of the source lying off the exact caustic fold position. While these will bias the estimation of the exact critical curve position (something not directly observed), it will not bias $\xi$, as this is sensitive only to uncorrelated shifts in the midpoint positions.
Appendix C The Effect of Individual Subhalos on Asymmetry
As can be seen in Figure 1, some realizations of asymmetry appear to be dominated by individual massive subhalos. This motivates exploring how the summary statistic $\xi$ is affected by the presence of individual massive subhalos. It can be shown that there is indeed a weak positive correlation between the most massive subhalo of a given realization and $\xi$. This correlation is intuitively expected, however, it remains subdominant due to considerable scatter in the relation. Furthermore, our statistical framework already accounts for the relative likelihood that an asymmetry comes from a single massive subhalo by averaging over many realizations. Therefore, this point is rendered a conceptual point rather than a true statistical effect in our method.
The more important effect with regards to our statistical method is the population level maximum subhalo mass. Our initial choice to simulate subhalos between $10^{6}-10^{10}M_{\odot}$ is consistent with previous works. However, recent work has explored the validity of this assumption. For example, Doppel et al. (2025) finds with TNG50 that truly observationally dark subhalos are most likely to have mass $\lesssim 10^{8.32}M_{\odot}$. This perhaps motivates reducing the upper bound of the mass range used for our simulation. Here, we vary the upper bounds of the simulation mass range for subhalos to study its effect on $\xi$.
Figure 15: Density distribution of the measured $\xi$ with $\delta_{xy}=0.01$" for various upper mass range bounds. The lower mass range bound is $10^{6}M_{\odot}$ in each case. Vertical lines indicate the median $\xi$ for each distribution.
We vary the upper mass range bound $M_{\rm upper}$, while keeping the lower mass range bound fixed. We sample 1000 realizations per $M_{\rm upper}$ bin, and only reduce $M_{\rm upper}$ in intervals of 0.5 (in log space). We do not consider increasing $M_{\rm upper}$ since subhalos with masses $10^{10}M_{\odot}$ are unlikely and are more likely to have an electromagnetic counterpart. Since we are interested in how this would change our results (which are derived from $\xi$), we include $\delta_{xy}=0.01$" for this test. Figure 15 shows these results. As can be seen, the distribution of $\xi$ is relatively constant across all $M_{\rm upper}$. There is a very small trend that can be observed where the median $\xi$ for each distribution slightly increases with $M_{\rm upper}$, however, it is statistically insignificant in this case since the fractional change in the median $\xi$ between each $M_{\rm upper}$ bin, relative to the $95\%$ confidence interval, is $\sim 2\%$. For these reasons, we conclude that changing $M_{\rm upper}$ would have a negligible effect on our results. Likewise, varying the lower mass range bound, while keeping the upper bound constant, also does not effect our results.
We conclude that the most massive subhalos in the simulation are a subdominant effect on $\xi$ that are already accounted for in our statistical model. In addition, properly accounting for the relative likelihood change from individual massive subhalos by varying the upper mass range bounds that we sample subhalos between has a very negligible effect on the results. We mention that in the future, greater care should be taken when determining the upper and lower mass range bounds in order to ensure that physically realistic subhalo masses are being used. This will be especially important once $\delta_{xy}$ becomes minimized for most systems.
Appendix D Testing the Convergence of the Posterior Distributions
We conduct a convergence test to determine if our results are dependent on the value of $N$, the number of realizations. To do so, we randomly discard one-third of the samples and recalculate the posterior with the remaining samples with the same acceptance criterion of the best 100. If the resampled posterior from a smaller $N$ remains mostly unchanged from before, then we can conclude that the posterior is converged. Because we are randomly discarding samples, there will be an associated scatter in the resampled posterior shape. To account for this, we resample 50 times and calculate the posterior medians. If the range in resampled posterior medians is low relative to the 68% confidence interval of the original posterior sample (from $N=20000$), then we conclude that the posterior is sufficiently converged relative to the uncertainty. When we divide these two quantities, we will refer to this as the relative median scatter, where a relative median scatter of 1.0 indicates that the resampled posterior median range is equal to the original 68% confidence interval. We note that this definition serves as an upper limit, as it can be susceptible to outliers.
We perform this test for the same mock arcs that we analyzed in Figure 7 of the main text. These results are shown in Figure 16, where the top and bottom panels correspond to the left and right columns of Figure 7. For clarity, we refer to these as the left and right mock arcs for this discussion. As can be seen, the resampled posteriors are relatively unchanged. We quantify this by dividing the resampled posterior median range (length of the red shaded region of Figure 16) with the original 68% confidence interval (length of the blue shaded region). We find that the left and right mock arcs exhibit a relative median scatter of 0.2 and 0.1, respectively. Because this scatter is much lower than the original posterior 68% confidence interval, we conclude that the original posteriors are stable and converged relative to the uncertainty. Repeating our statistical method with larger values of $N$ will only improve the precision of the results, and we encourage doing so in future works, provided the appropriate computational resources.
Figure 16: Convergence test posterior results for the two perpendicular mock arcs shown in Figure 7 of the main text. The top and bottom panels correspond to the same arcs as those in the left and right columns of Figure 7, respectively. The black lines indicate the original posterior, the same as the dashed black line in Figure 7. The red lines show resampled posteriors after discarding one-third of the original posterior samples. The blue dashed line and shaded region indicate the original posterior medians and 68% confidence interval. The red shaded region indicates the full posterior median range from the resampled posteriors.